OpenStudy (anonymous):

find the solution set of x? 1/2x-3>o can anyone help me?

OpenStudy (anonymous):

\[\frac{1}{2}x-3>0\] \[\frac{1}{2}x>3\] \[x>\frac{3}{2}\]